OpenStudy (anonymous):

1.)9.35m+3.2m+7.81m ?? 2.)8.32m-2.21m ?? 3.)8.723cm x 2.36cm ??? 4.)0.7722m/0.11 cm ??

OpenStudy (anonymous):

1.)9.35m+3.2m+7.81m ?? 2.)8.32m-2.21m ?? 3.)8.723cm x 2.36cm ??? 4.)0.7722m/0.11 cm ?? plss help me..

OpenStudy (anonymous):

is that calculation of units? if so than: 9,35m+3,2m+7,81m = 20,36m 8,32m-2,21m= 6,11m 8,723 cm x 2,36 cm= 20,58628 cm^2 0,7722 m / 0,11*10-2m= 702 m

OpenStudy (anonymous):

oW.. tnx for helping me Kryten..

OpenStudy (anonymous):

not a problem

OpenStudy (anonymous):

in number is divide.. 0.7722m divided by 0.11 cm

OpenStudy (anonymous):

in number 4...

OpenStudy (anonymous):

use round off...

OpenStudy (anonymous):

yes and 0,11 cm = 0,11 *10^-2 m because 1 m = 100 cm

OpenStudy (anonymous):

ha............ ok ......... thanks.

OpenStudy (anonymous):

that should be right